# Stability for Strichartz inequalities: Existence of minimizers

### v1

@@ Abstract.
We study the quantitative stability associated to adjoint Fourier restriction inequality, focusing on the paraboloid and two-dimension sphere cases. We show that these Strichartz-stability inequalities admit minimizers attaining their sharp constants, on the condition that these sharp constants are strictly smaller than the corresponding spectral-gap constants. Furthermore, for the two-dimension sphere case, we obtain the existence of minimizers.
@@
This work was supported in part by the National Key R&D Program of China [Grant No. 2023YFC3007303], National Natural Science Foundation of China [Grant Nos. 12071052 & 12271501 & 12471232 & 12501130], and China Postdoctoral Science Foundation [Grant Nos. GZB20230812 & 2024M753436].
@@ 1. Introduction
The study of Fourier restriction phenomena plays a central role in harmonic analysis over the past decades. Roughly speaking, the well-known Paraboloid/Sphere Fourier Restriction Conjecture due to Stein asks the following question: for some given exponents $(p,q)$, does the associated $L^{p}\to L^{q}$ adjoint Fourier restriction inequalities hold? For the case $p=2$, this question was fully solved by Tomas-Stein [Tom75]. In this paper, we focus on this Tomas-Stein inequalities for the paraboloid and sphere, which are also known as Strichartz inequalities [Str77] since the space-time Fourier transform of PDE-solutions are supported on these surfaces.
@@
There are natural questions whether these Strichartz inequalities (1) admit maximizers, and moreover, whether these Strichartz inequalities are stable in the sense that this Strichartz deficit functional $\delta_{S}(f)$ can control this distance from the maximizer manifold $\mathrm{dist}(f,\mathcal{M})$. In fact, the maximizers are conjectured to be Gaussian functions for the paraboloid case, and conjectured to be constant functions for the sphere case; while these two conjectures are only confirmed for the paraboloid with $d=\{1,2\}$ by Foschi [Fos07] and Hundertmark-Zharnitsky [HZ06], then confirmed for the sphere with $d=2$ by Foschi [Fos15]. For more progress on this maximizer problems we refer to the survey papers [FOeS17, NOeST23, OeS24]. Next, for the quantitative stability of Strichartz inequalities [^1] Indeed, the precompactness of maximizing sequences can give a qualitative stability result, further details on this aspect are referred to [BOeSQ20, DY24, CS12, FLS16, Kun03, Ram12, Sha09, Sha16] and the references therein., Duyckaerts-Merle-Roudenko[DMR11] have shown the coercivity of a relevant quadratic form for the paraboloid with $d=\{1,2\}$; and recently, Gonçalves-Negro[GN22] have shown the following exactly stability estimates for the paraboloid and sphere cases [^2] They have also studied the wave equation case, which essentially corresponds to the cone case. See also the papers [Gon19a, Gon19b, Neg23a, Neg23b] for similar quantitative stability on Fourier restriction type inequalities.: there exist $\eta\in(0,1)$ and $C_{\sharp\sharp}>0$ such that
@@
on the condition that $\mathrm{dist}(f,\mathcal{M})<\eta\|f\|_{L^{2}(S)}$; and this condition can be removed if the corresponding maximizer conjecture is confirmed. Similarly, we call a nonzero function $f_{0}\in L^{2}(S)$ a minimizer for $C_{\sharp\sharp}$ if it can make the inequality (2) an equality.
However, the results in [GN22] do not give too much information on the stability constants $C_{\sharp\sharp}$. On the one hand, what is the optimal value of this stability constant $C_{\sharp\sharp}$; on the other hand, whether this value $C_{\sharp\sharp}$ can be attained for a minimizer. As a comparison, we recall the corresponding research progress on the Sobolev inequality: the maximizers are exactly Aubin-Talenti manifold according to the celebrated works [Aub76, Lie83, Ros71, Tal76], then the Sobolev-stability questions are raised by Brézis-Lieb [BL85] and the corresponding Sobolev stability estimate is established by Bianchi-Egnell [BE91], recently the explicit stability constant is investigated by Dolbeault-Esteban-Figalli-Frank-Loss [DEF+25] whose result is dimensionally sharp, and the existence of minimizers for stability constant is obtained by König [K2̈5]. For further details on this aspect, we refer to the survey papers [DEF+24, Fra24]. Here we should emphasize that in the Sobolev setting, by the work of [K2̈5, Section 1.1], there are two crucial phenomena (two-peak and spectral-gap) which both play a role and thus lead to two compactness thresholds.
@@
  • for the existence of minimizers for $C_{\sharp\sharp}$, we discover an interesting fact that the sphere and paraboloid cases exhibit quite different behaviors: we obtain the existence of minimizers for the two-dimension sphere, but only conditional existence of minimizers for the paraboloid.
Our main results are summarized as follows. Let $C_{*}$ and $C_{**}$ denote the sharp stability constants in (2) for the paraboloid and the two-dimension sphere situation respectively, see also the precise expressions in the following (5) and (7). Then we have
@@
We indeed have shown that $C_{*}\leq\widetilde{C}_{d}$ must hold, and in fact, the equality $C_{*}=\widetilde{C}_{d}$ will imply the noncompactness [^3] Note that this noncompactness does not imply the nonexistence of minimizers. of minimizing sequences for $C_{*}$. We mention that, for this condition (3), similar strict inequalities hold for the Sobolev stability inequality [K2̈3] and planar isoperimetric stability inequality [BCH17], while their methods do not seem to work in our setting.
In the following two subsections, we present some precise terminologies on the paraboloid case and two-dimension sphere case respectively. Meanwhile, we explain the main difficulties of our problem and the main methods of our proof.
@@ 1.1. Paraboloid adjoint Fourier restriction
thus one can directly see that $C_{*}\leq C_{SG}$. In this paper, based on [GN22, Proof of Theorem 1.2], we can further investigate some monotonicity to show the following Proposition 1.6, which guarantees that the spectral-gap obstacle cannot appear due to the the condition (3).
@@
In summary, if there holds the strict inequality $C_{*}<\min\{C_{SG},C_{TP}\}$, then the spectral-gap and two-peak both cannot happen, and thus it is expected that the existence of minimizers can be established. In fact, by some direct computation, we can prove that the two-peak is vanished in our Strichartz setting, which comes from the following result.
@@
1.2. Two-dimension sphere adjoint Fourier restriction
@@
$\delta_{*}(f):=\mathbf{M}_{d}^{2}\|f\|_{L^{2}(\mathbb{S}^{d-1})}-\left\|\widehat{f\sigma}\right\|_{L^{q}(\mathbb{R}^{d})}^{2},\quad\mathrm{dist}(f,\mathcal{C}):=\inf_{g\in\mathcal{G}}\left\|f-g\right\|_{L^{2}(\mathbb{S}^{d-1})}.$
@@
(7) $\delta_{*}(f)\geq C_{**}\mathrm{dist}(f,\mathcal{C})^{2},\quad C_{**}:=\inf_{0\neq f\in L^{2}}\frac{\delta_{*}(f)}{\mathrm{dist}(f,\mathcal{G})^{2}}.$
@@
$h\perp 1\quad\Rightarrow\quad\delta(1+\varepsilon h)=C_{SG}\|h\|_{L^{2}(\mathbb{R}^{d})}^{2}\varepsilon^{2}+o(\varepsilon^{2}),$
thus one can directly see that $C_{*}\leq C_{SG}$. In this paper, by investigating some monotonicity in the work of Gonçalves-Negro [GN22, Proof of Theorem 1.3] and combining some spectral perturbative calculation due to König [K2̈3], we can show the following Proposition 1.11, which directly guarantees that the spectral-gap obstacle cannot appear.
@@
Finally, it is obvious that the two-peak phenomenon is vanished since $C_{SG*}<C_{TP*}$, and moreover $C_{**}<\min\{C_{SG*},C_{TP*}\}$. Therefore the aforementioned two obstacles are both disappeared, and we thus can obtain the desired unconditional existence of minimizers Theorem 1.1.
@@ 1.3. Outline of the paper
We show the paraboloid result Theorem 1.3 in Section 2, and show the two-sphere result Theorem 1.1 in Section 3. In each section, we proceeds three subsections to investigate the spectral-gap, two-peak, and existence of minimizers respectively.
@@ 2.1. Spectral-gap
over all $T_{G}\mathcal{G}^{\perp}$ is the same as the minima over $T_{G}\mathcal{G}^{\perp}\cap L_{rad}^{2}(\mathbb{R}^{d})$; moreover for any function $f\in L_{rad}^{2}(\mathbb{R}^{d})$ with expression
@@
there holds
@@ 2.3. Conditional compactness of minimizing sequences
The proof strategy is very similar to the proof of [K2̈5, Theorem 1.2]. The arguments are pretty long, but standard, as shown in [K2̈5, Section 4]. Here for simplicity, we briefly summary the proof steps and omit the details.
@@
3. Existence of minimizers: two-dimension sphere
@@ 3.1. Spectral-gap
where $Y_{k}$ is an orthogonal set of real spherical harmonics with $\|Y_{k}\|_{L^{2}(\mathbb{S}^{2})}=1$, there holds
@@
Then, similar to Step 1, by some direct computation [^6] One can also use some software such as Mathematica to check these integrals., on can obtain
@@ 3.3. Compactness of minimizing sequences
In this subsection, based on the spectral-gap result Proposition 1.11 and two-peak vanishing result Proposition 1.12, we show the desired existence of minimizers for the two-dimension sphere stability constant $C_{**}$.
@@
The strategy is similar to the [K2̈5, Section 4, Proof of Theorem 1.2] and the Proof of Theorem 1.3 in the previous Section 2.3. For simplicity, we only show the main ideas and omit the details. In this two-dimension sphere case, we can obtain the unconditional result since we can establish the strict inequality in Proposition 1.11. Therefore, by following the outline in the Proof of Theorem 1.3, it is only need to check the fact

### v2

@@ Abstract.
We study the quantitative stability associated with the adjoint Fourier restriction inequality, focusing on the paraboloid and two-dimensional sphere cases. We show that these Strichartz-stability inequalities admit minimizers attaining their sharp constants, provided that these sharp constants are strictly smaller than the corresponding spectral-gap constants. Furthermore, for the two-dimensional sphere case, we obtain the existence of minimizers.
@@
This work was supported in part by the National Key R&D Program of China [Grant No. 2023YFC3007303], National Natural Science Foundation of China [Grant Nos. 12071052 & 12271501 & 12471232 & 12501130], and China Postdoctoral Science Foundation [Grant Nos. GZB20230812 & 2024M753436 & YJC20250892].
@@ 1. Introduction
The study of Fourier restriction phenomena plays a central role in harmonic analysis over the past decades. Roughly speaking, the well-known Paraboloid/Sphere Fourier Restriction Conjecture due to Stein asks the following question: for some given exponents $(p,q)$, do the associated $L^{p}\to L^{q}$ adjoint Fourier restriction inequalities hold? For the case $p=2$, this question was fully solved by Tomas-Stein [Tom75]. In this paper, we focus on these Tomas-Stein inequalities for the paraboloid and sphere, which are also known as Strichartz inequalities [Str77] since the space-time Fourier transform of PDE-solutions are supported on these surfaces.
@@
There are natural questions whether these Strichartz inequalities (1) admit maximizers, and moreover, whether these Strichartz inequalities are stable in the sense that this Strichartz deficit functional $\delta_{S}(f)$ can control this distance from the maximizer manifold $\mathrm{dist}(f,\mathcal{M})$. In fact, the maximizers are conjectured to be Gaussian functions for the paraboloid case, and conjectured to be constant functions for the sphere case; while these two conjectures are only confirmed for the paraboloid with $d=\{1,2\}$ by Foschi [Fos07] and Hundertmark-Zharnitsky [HZ06], then confirmed for the two-dimensional sphere $\mathbb{S}^{2}\subset\mathbb{R}^{3}$ by Foschi [Fos15]. For more progress on this maximizer problems we refer to the survey papers [FOeS17, NOeST23, OeS24]. Next, for the quantitative stability of Strichartz inequalities [^1] Indeed, the precompactness of maximizing sequences can give a qualitative stability result, further details on this aspect are referred to [BOeSQ20, DY24, CS12, FLS16, Kun03, Ram12, Sha09, Sha16] and the references therein., Duyckaerts-Merle-Roudenko[DMR11] have shown the coercivity of a relevant quadratic form for the paraboloid with $d=\{1,2\}$; and recently, Gonçalves-Negro[GN22] have shown the following quantitative stability estimates for the paraboloid and sphere cases [^2] They have also studied the wave equation case, which essentially corresponds to the cone case. See also the papers [Gon19a, Gon19b, Neg23a, Neg23b] for similar quantitative stability on Fourier restriction type inequalities. In a related direction, see [CNOeS24] for weighted-perturbation stability of maximizers in spherical extension inequalities.: there exist $\eta\in(0,1)$ and $C_{\sharp\sharp}>0$ such that
@@
under the condition that $\mathrm{dist}(f,\mathcal{M})<\eta\|f\|_{L^{2}(S)}$; and this condition can be removed if the corresponding maximizer conjecture is confirmed. Similarly, we call a nonzero function $f_{0}\in L^{2}(S)$ a minimizer for $C_{\sharp\sharp}$ if it can make the inequality (2) an equality.
However, the results in [GN22] do not give too much information on the stability constants $C_{\sharp\sharp}$. On the one hand, what is the optimal value of this stability constant $C_{\sharp\sharp}$; on the other hand, whether this value $C_{\sharp\sharp}$ can be attained for a minimizer. As a comparison, we recall the corresponding research progress on the Sobolev inequality: the maximizers are exactly Aubin-Talenti manifold according to the celebrated works [Aub76, Lie83, Ros71, Tal76], then the Sobolev-stability questions are raised by Brézis-Lieb [BL85] and the corresponding Sobolev stability estimate is established by Bianchi-Egnell [BE91], recently the explicit stability constant is investigated by Dolbeault-Esteban-Figalli-Frank-Loss [DEF+25] whose result is dimensionally sharp, and the existence of minimizers for stability constant is obtained by König [K2̈5]. For further details on this aspect, we refer to the survey papers [DEF+24, Fra24] and the references therein, see also [Chr14, Chr17, CLT23, CLT24, DSW25, FMP08, Fus15, FZ22] for quantitative stability in different settings. Here we emphasize that in the Sobolev setting, by the work of [K2̈5, Section 1.1], there are two crucial phenomena (two-peak and spectral-gap) which both play a role and thus lead to two compactness thresholds.
@@
  • for the existence of minimizers for $C_{\sharp\sharp}$, we discover an interesting fact that the sphere and paraboloid cases exhibit quite different behaviors: we obtain the existence of minimizers for the two-dimensional sphere, but only conditional existence of minimizers for the paraboloid.
Our main results are summarized as follows. Let $C_{*}$ and $C_{**}$ denote the sharp stability constants in (2) for the paraboloid and the two-dimensional sphere situation respectively, see also the precise expressions in the following (5) and (7). Then we have
@@
We indeed have shown that $C_{*}\leq\widetilde{C}_{d}$ must hold, and in fact, the equality $C_{*}=\widetilde{C}_{d}$ will imply the noncompactness [^3] Note that this noncompactness does not imply the nonexistence of minimizers. of minimizing sequences for $C_{*}$. We mention that, for this condition (3), similar strict inequalities hold for the quantitative stability of Sobolev inequality [K2̈3] and planar isoperimetric inequality [BCH17], while their methods do not seem to work in our setting.
In the following two subsections, we present some precise terminologies on the paraboloid case and two-dimensional sphere case respectively. Meanwhile, we explain the main difficulties of our problem and the main methods of our proof.
@@ 1.1. Paraboloid adjoint Fourier restriction
thus one can directly see that $C_{*}\leq C_{SG}$. Indeed, Gonçalves-Negro [GN22, Proof of Theorem 1.2] has shown that $C_{SG}>0$. In this paper, we can further investigate some monotonicity to show the following Proposition 1.6, which guarantees that the spectral-gap obstacle cannot appear due to the the condition (3).
@@
In summary, if there holds the strict inequality $C_{*}<\min\{C_{SG},C_{TP}\}$, then the spectral-gap and two-peak both cannot happen, and thus it is expected that the existence of minimizers can be established. In fact, by some direct computation, we can prove that the two-peak vanishes in our Strichartz setting, which comes from the following result.
@@
1.2. Two-dimensional sphere adjoint Fourier restriction
@@
$\delta_{*}(f):=\mathbf{M}_{d}^{2}\|f\|_{L^{2}(\mathbb{S}^{d-1})}-\left\|\widehat{f\sigma}\right\|_{L^{q}(\mathbb{R}^{d})}^{2},\quad\mathrm{dist}(f,\mathcal{C}):=\inf_{g\in\mathcal{C}}\left\|f-g\right\|_{L^{2}(\mathbb{S}^{d-1})}.$
@@
(7) $\delta_{*}(f)\geq C_{**}\mathrm{dist}(f,\mathcal{C})^{2},\quad C_{**}:=\inf_{0\neq f\in L^{2}}\frac{\delta_{*}(f)}{\mathrm{dist}(f,\mathcal{C})^{2}}.$
@@
$h\perp 1\quad\Rightarrow\quad\delta(1+\varepsilon h)=C_{SG*}\|h\|_{L^{2}(\mathbb{R}^{d})}^{2}\varepsilon^{2}+o(\varepsilon^{2}),$
thus one can see that $C_{**}\leq C_{SG*}$. Indeed, Gonçalves-Negro [GN22, Proof of Theorem 1.3] has shown that $C_{SG*}>0$. In this paper, by investigating some monotonicity in the work of Gonçalves-Negro and combining some spectral perturbative calculation due to König [K2̈3], we can show the following Proposition 1.11, which directly guarantees that the spectral-gap obstacle cannot appear.
@@
Finally, it is obvious that the two-peak phenomenon vanishes since $C_{SG*}<C_{TP*}$, and moreover $C_{**}<\min\{C_{SG*},C_{TP*}\}$. Therefore the aforementioned two obstacles are eliminated, and we thus can obtain the desired unconditional existence of minimizers Theorem 1.1.
@@ 1.3. Outline of the paper
We show the paraboloid result Theorem 1.3 in Section 2, and show the two-sphere result Theorem 1.1 in Section 3. In each section, we proceed three subsections to investigate the spectral-gap, two-peak, and existence of minimizers respectively.
@@ 2.1. Spectral-gap
over all $T_{G}\mathcal{G}^{\perp}$ is the same as the minima over $T_{G}\mathcal{G}^{\perp}\cap L_{rad}^{2}(\mathbb{R}^{d})$; moreover for any function $f\in L_{rad}^{2}(\mathbb{R}^{d})$ with the following Laguerre polynomials expression
@@
where $L_{m}^{d/2-1}$ is the standard Laguerre polynomial, there holds
@@ 2.3. Conditional compactness of minimizing sequences
The proof strategy is very similar to the proof of [K2̈5, Theorem 1.2]. The arguments are pretty long, but standard, as shown in [K2̈5, Section 4]. Here for simplicity, we briefly summarize the proof steps and omit the details.
@@
3. Existence of minimizers: two-dimensional sphere
@@ 3.1. Spectral-gap
where $Y_{k}$ is the standard real spherical harmonics with $\|Y_{k}\|_{L^{2}(\mathbb{S}^{2})}=1$, there holds
@@
Then, similar to Step 1, by some direct computation [^6] One can also use some software such as Mathematica to check these integrals., one can obtain
@@ 3.3. Compactness of minimizing sequences
In this subsection, based on the spectral-gap result Proposition 1.11 and two-peak vanishing result Proposition 1.12, we show the desired existence of minimizers for the two-dimensional sphere stability constant $C_{**}$.
@@
The strategy is similar to the [K2̈5, Section 4, Proof of Theorem 1.2] and the Proof of Theorem 1.3 in the previous Section 2.3. For simplicity, we only show the main ideas and omit the details. In this two-dimensional sphere case, we can obtain the unconditional result since we can establish the strict inequality in Proposition 1.11. Therefore, by following the outline in the Proof of Theorem 1.3, it remains only to check that